# StrCGAN: A Generative Framework for Stellar Image Restoration

### v2

@@ 4.2 Development of StrCGAN
5 Discussion
This section presents a qualitative exploration of the factors underlying StrCGAN’s apparent outperformance compared to previous models, an analysis of its architectural evolution, and the key lessons learned from the experiments. The discussion is grounded in the visual and conceptual outcomes observed during experimentation with the MobilTelesco dataset [1], offering insights into the design choices that enabled StrCGAN’s success in enhancing low-resolution astrophotographic images.
5.1 Qualitative Outperformance of StrCGAN
StrCGAN’s improved performance, as inferred from qualitative assessments, can be attributed to several key factors that address the limitations of prior models. Unlike DCGAN [31], BigGAN [6], StyleGAN [19], CGAN [27], Pix2Pix [16], DDPM [13], and DDIM [37]—which frequently produced noisy or artifact-ridden outputs with little to no discernible stellar morphology—StrCGAN consistently delivered visually coherent reconstructions. For example, inference results on the test split, particularly for objects such as Bellatrix and Aldebaran, revealed detailed star shapes and features that were absent in outputs from other models, even after extensive hyperparameter tuning and data augmentation. This suggests that StrCGAN’s ability to capture and preserve astrophysical structures stems from its tailored approach to challenges specific to the dataset, such as atmospheric distortion and low SNR nature [22, 10].
The integration of 3D convolutional processing likely played a pivotal role, enabling StrCGAN to model volumetric data in a way that 2D-based models such as Pix2Pix and StyleGAN could not. This volumetric representation allowed for a more accurate depiction of celestial depth, which is critical in astronomy where both surface topology and surface appearance are essential [36, 9]. Additionally, the multi-spectral fusion of Optical and NIR data from All Sky Surveys provided richer contextual information, enabling StrCGAN to align features across spectral domains more effectively than single-domain baselines such as SinGAN [35] and ConSinGAN [12]. The qualitative leap in inference quality, even on unpaired Jupiter images, underscores StrCGAN’s robustness and adaptability, suggesting that it overcame the feature-deficiency issues that hindered previous approaches [30].
5.2 Architectural Differences from Previous Versions
StrCGAN represents a significant evolution from its predecessors, including the initial CycleGAN-inspired framework [46] and intermediate iterations such as CMRA_GAN and PiSGAN. These architectural differences are central to its improved performance and reflect a series of iterative refinements based on experimental feedback. Unlike the original CycleGAN, which relied solely on 2D image-to-image translation with a U-Net-like generator and PatchGAN discriminator, StrCGAN introduces 3D convolutional layers (adaptable via nn.Conv3d), enabling it to process volumetric data and capture depth information critical for astrophotographic enhancement. This shift from 2D to 3D processing marks a departure from earlier versions, which struggled with flat representations that failed to preserve stellar topology.
The addition of a multi-resolution attention module, implemented as nn.Sequential(nn.Conv2d(256, 1, 1), nn.Sigmoid()), further distinguishes StrCGAN from previous architectures. This module allows the model to emphasize fine details across different scales, a capability absent in earlier designs such as LumenGAN, which lacked adaptive focus and often blurred critical stellar features [41]. The multi-spectral fusion layer, which concatenates Optical and NIR features, represents another advancement over AstroVISTA, which processed single spectral bands and thus missed the cross-domain alignment that enhances StrCGAN’s fidelity. Moreover, the astrophysical regularization—custom loss terms penalizing deviations from stellar morphology—adds a domain-specific constraint absent in CMRA_GAN, where outputs sometimes deviated into non-astronomical artifacts [9]. Collectively, these architectural enhancements enabled StrCGAN to produce visually superior reconstructions compared to the noisy or distorted outputs of its predecessors.
5.3 Lessons Learned from PiSGAN
Our experiments with PiSGAN, a physics-informed GAN based on the stacked GAN (SGAN)[15] framework, introduced domain-specific constraints (Poisson noise, SNR) into the loss function to enforce astrophysical realism. Its U-Net generator (7 encoders, 6 decoders) processed high-resolution inputs (3×1536×2048), while stacked discriminators with spectral normalization stabilized training and evaluated both image fidelity and physical plausibility.
Despite these advances, PiSGAN underperformed. Outputs contained artifacts and unstable losses, and FID (309.4) remained worse than baseline GANs and diffusion models. The 2D U-Net struggled with volumetric and spectral structure, and reliance on physics constraints without unpaired training or morphological regularization limited generalization. Early termination at 49 epochs further prevented adaptation.
Key takeaways were: (i) volumetric and spectral complexity requires 3D and multi-resolution attention mechanisms; (ii) unpaired translation is essential to overcome scarce paired data; and (iii) astrophysical regularization must directly enforce morphology. These lessons directly motivated StrCGAN’s architecture and training strategy, which achieved qualitatively superior reconstructions.
6 Conclusion and Future Work
@@
In this work, we introduced StrCGAN, a physics-informed conditional GAN for enhancing low-resolution astrophotographic images using limited MobilTelesco data. By combining 3D convolutions, multi-spectral fusion, and astrophysical regularization, StrCGAN preserved stellar morphology and fine structures, outperforming prior GAN and diffusion models in qualitative reconstruction. However, its generalization remains constrained by the scarcity of ground-truth references and the high computational demands of volumetric processing.
@@
Future work will expand the dataset with additional multi-spectral and time-lapse observations, leverage unpaired data from modern all-sky surveys, and explore scalable architectures integrating transformer-based attention and temporal modeling. These extensions aim to improve quantitative performance, enable dynamic stellar reconstruction, and embed stronger physical priors, moving toward models that are both visually and scientifically informative.
Acknowledgments
I would like to thank Herr Silas Janke from the Department of Physics and Astronomy, Heidelberg University (ORCID: 0009-0000-4192-5390), for his invaluable support as a data curator. His assistance in setting up the ground-truth datasets for the all-sky survey optical and near-infrared data catalogs was crucial to this work.

### v3

@@ 6.3 Future Directions
7 Conclusion
@@
In this work, we introduced StrCGAN, a physics-informed conditional GAN for enhancing low-resolution astrophotographic images using limited MobilTelesco data. By combining multi-spectral fusion, and astrophysical regularization, StrCGAN attempted to boost pixel morphology but underperformed to exisiting GAN and diffusion models in quantitative evaluation. However, its visual reconstruction was better compared to the benchmarks on those other models. Generalization of its used remains constrained by the scarcity of ground-truth references and the high computational demands of extended processing.
@@
Future work will expand the training set with additional multi-spectral and time-lapse observations, leverage unpaired data from modern all-sky surveys, and explore scalable architectures integrating transformer-based attention and temporal modeling. These extensions aim to improve quantitative performance, enable dynamic stellar reconstruction, and embed stronger physical priors, moving toward models that are both visually and scientifically improved.
Figure 6: Enhanced astronomical images produced by our final STRCGAN model. All outputs are direct model predictions (inference mode) from real mobile-phone captures.
Figure 7: Visual comparison conventional Generative models trained over MobilTelesco, highlighting the lack of features and morphology capturing capability.
Figure 8: Results from our previous PISGAN baseline using different PSF modeling strategies. Top three rows: physically-motivated Custom PSF. Bottom two rows: Random PSF. While PISGAN already outperforms conventional baselines, our final STRCGAN (Fig. 6) shows significantly superior detail preservation and artifact suppression.